OpenStudy (anonymous):

How do I simplify: 2√7 + √5 / 3√7 - 2√5? The answer is 52 + 7√35 / 43.

OpenStudy (mimi_x3):

is it: \[\large \frac{2\sqrt{7}+5\sqrt{5}}{3\sqrt{7}-2\sqrt{5}} \] ?

OpenStudy (anonymous):

yup! I got it! thanks anyway! :P

OpenStudy (mimi_x3):

lool, im way too slow xD

OpenStudy (anonymous):

but

OpenStudy (anonymous):

if you don't mind: How do i simplify: √2 + √5 / √6 - √10?

OpenStudy (anonymous):

I got √112 / -√64 but it's not the right answer.

OpenStudy (anonymous):

The answer is supposed to be: - 5√2 + 2√3 + 2√5 + √30 / 4